# GroundingBooth: Grounding Text-to-Image Customization

### v1

@@ 4.3 Model Training
Based on the modified masked cross-attention, the layout-guided training pipeline is formulated as:
$\mathcal{L}(\theta)=\mathbb{E}_{z,l_{obj},y,x,\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(z_{t},t,l_{obj},c_{t}(y),c_{i}(x)\right)\right\|_{2}^{2}\right],$ (9)
where $z$ is the latent space reference image, $l_{obj}$ is the input layout control, $y$ is the text input, $x$ is the input reference object,$\epsilon$ is the noise term, $t$ is the time step, $\epsilon_{\theta}$ is the denoising U-Net, $z_{t}$ is the noisy version of $z$ at time $t$, $c_{t}$ is the CLIP text encoder and $c_{i}$ is the DINOv2 image encoder. During training, we keep the text encoder and DINOv2 image encoder frozen and merely fine-tune the self-attention layers, the gated self-attention layers, the masked cross-attention layers, and the MLP after the DINOv2 image encoder.
@@ 4.4 Model Inference
Although our model is trained on single-reference object data, it has the extended ability to achieve multi-reference object customization. As shown in Fig. 2, in the inference stage, assume we have $n$ reference objects, the reference object and paired layout information are concatenated as the grounded reference image embeddings. In each transformer block, the masked cross-attention layer is copied for $n$ times, and each ID token and its paired layout information are injected into each masked cross-attention layer respectively. Owning to the restriction of the masked cross-attention module, both the independence of generation between foreground objects and background text entities and the generation within foreground objects are properly maintained. In this way, our model allows for grounded generation of any-number reference object entities and any-number text entities, achieving flexible and precise layout control for both multiple reference foreground objects and text-driven background.
@@ 5.1 Dataset
We conduct our experiments on datasets from two sources. To train the identity preservation ability of the model, we adopt both multiview datasets and single-image datasets. For multi-view data, we use MVImgNet Yu et al. (2023), which contains 6.5 million frames from 219,188 videos crossing objects from 238 classes, with fine-grained annotations of object masks. In the data processing stage, following AnyDoor Chen et al. (2023b), for each object, we use randomly selected different frames of the video clips. For the first frame, we apply the mask on the image and get the foreground object, and for another frame, we use the bounding box of the object as the grounding information and use this frame as the training ground truth. For single-image data, we use LVIS Gupta et al. (2019), a well-known dataset for fine-grained large vocabulary instance segmentation, including 118,287 images of 1203 categories. For one image, we select the objects with top-10 bounding box area for training, in which we use the object with the largest bounding box as the reference object while the other objects as used as background.
@@ 5.2 Evaluation Metrics
We calculate the CLIP-I Radford et al. (2021) score and DINO Caron et al. (2021) score to evaluate the identity preservation performance of the model and use CLIP-T Radford et al. (2021) score to evaluate the text alignment performance of text-to-image generation.
Figure 5: Visual comparison with existing methods based on DreamBench Objects.
Figure 6: Multi-concept Generation based on DreamBench objects.
5.3 Comparison with Previous Works
We compare our work with previous state-of-the-art works, which are shown in Fig. 5 and Table 1. Compared with previous encoder-based subject-driven text-to-image customization methods like BLIP-Diffusion Li et al. (2024), ELITE Wei et al. (2023), $\lambda$-eclipse Patel et al. (2024) and MLLM-based method Kosmos-G Pan et al. (2023), our method maintains accurate identity of the reference objects, and achieve strong grounding ability, enabling flexible layout control of the reference objects. Methods like AnyDoor Chen et al. (2023b) can only deal with a given background, which lacks the ability of text-to-image generation. Although previous grounded text-to-image generation methods like GLIGEN Li et al. (2023) is able to achieve precise layout control, it cannot maintain the identity of the objects. CustomNet Yuan et al. (2023) achieves flexible pose control, while it highly relies on the pretrained model from Zero123 Liu et al. (2023), restricting its generation resolution to be $256\times 256$. Our method shows better performance in layout alignment, foreground object identity preservation and background text alignment.
5.4 Multi-Object Foreground-Background Joint Control
We show qualitative results of multi-object and multi-background text entities in Fig. 6. Results show that our method achieves joint foreground-background control with text alignment and identity preservation of foreground objects. Even when the bounding box of the foreground objects has a large overlap with the background text entities, the model can distinguish subject-driven foreground generation from text-driven background generation, effectively avoiding context blending and remaining scene hallucination.
Figure 7: Reference-guided layout-to-image synthesis result on COCO validation set. Note that LayoutDiffusion (Zheng et al., 2023a) is only conducted on COCO dataset with filtered annotations, so some of its results are not available.
5.5 Complex background Evaluation
We also evaluate our model’s performance of both identity preservation and grounding on COCO validation set, shown in Table 2 and Fig. 7 respectively. We use the object with the largest area as the reference object, and the remaining text entities as background entities. Results show that even in complex scenes where there are various background objects, providing their bounding boxes, our model generates results with precise layout, which not only shows obvious identity preservation of the reference object but also accurate layout-alignment and text-alignment, achieving flexible layout customized text-to-image generation. Compared with previous layout-to-image generation methods, our model has competitive accuracy in grounding and significant improvement in identity preservation evaluation.
Table 1: Comparison with previous methods on Dreambench.
@@ 5.5 Complex background Evaluation
SD V1.4 [Rombach et al. (2022a)] 0.3122 0.8413 0.6587
BLIP-Diffusion [Li et al. (2024)] 0.2824 0.8894 0.7625
ELITE [Wei et al. (2023)] 0.2461 0.8936 0.7557
Kosmos-G [Pan et al. (2023)] 0.2864 0.8452 0.6933
lambda-eclipse [Patel et al. (2024)] 0.2767 0.8973 0.7934
AnyDoor [Chen et al. (2023b)] 0.2430 0.9062 0.7928
GLIGEN [Li et al. (2023)] 0.2898 0.8520 0.6890
CustomNet [Yuan et al. (2023)] 0.2815 0.9090 0.7526
@@
Table 2: Evaluation on MS-COCO validation set: In this setting we finetune our model on COCO training set, and compare with previous methods that only train on COCO. To make fair comparison, following LayoutDiffusion, we compare results on filtered COCO validation set with 3097 images.
CLIP-T ↑ CLIP-I ↑ DINO-I ↑
LAMA[ Li et al. (2021)] 0.2507 0.8441 0.7330
LayoutDiffusion[ Zheng et al. (2023b)] 0.2738 0.8655 0.8033
GLIGEN[ Li et al. (2023)] 0.2899 0.8688 0.7792
Ours 0.2946 0.9078 0.8560
5.6 Abalation Study
We conduct the ablation study in two parts: the first part is for our cross-attention control, and the second part is for the grounding module. The results are shown in Table 4 and Table 4. From the results, we can conclude that both the grounding module and the masked cross-attention module play a vital role in grounding the reference object and preventing the information leakage of the reference object.
@@ 5.6 Abalation Study
CLIP-T ↑ CLIP-I ↑ DINO-I ↑
W/o Grounding Module 0.2796 0.8605 0.7740
W/o Masked Cross-attention 0.2884 0.8707 0.7970
Full 0.2945 0.9008 0.8555
6 Conclusion

### v2

@@ 3.3 Model Training
During training, for each image, we input only one subject image and its bounding box to the model, along with several text entities with their corresponding bounding boxes. The number of entities per training image is limited to 10 and we drop the rest ones. For a portion of training cases, the input may not contain subject image or text entities. We keep the text encoder and DINOv2 image encoder frozen and merely fine-tune the gated self-attention layers, the masked cross-attention layers, and the multi-layer perceptron after the DINOv2 image encoder.
@@ 3.4 Model Inference
Although our model is trained on single-subject data, it can be seamlessly extended to achieve multi-subject customization without retraining. As shown in Fig. 2, in the inference stage, assume we have $n$ reference objects, the reference object and paired layout information are concatenated as the grounded reference image embeddings. In each transformer block, the masked cross-attention layer will be reused for $n$ times, and each ID token and its paired layout information are injected into each masked cross-attention layer respectively. As we analyzed in section 3.2, our masked cross-attention ensures the independence of the generation of each subject, preventing potential false blending of visual concepts, e.g., the unnatural blending of two objects in the overlapping regions. It also guarantees an accurate layout control on all the subjects.
@@ 4.1 Dataset
The training data of our experiments are from both multi-view datasets and single-image datasets. For multi-view data, we use MVImgNet (Yu et al., 2023), which contains 6.5 million frames from 219,188 videos across 238 object categories, with fine-grained annotations of object masks. In the data processing stage of MVImgNet, following AnyDoor (Chen et al., 2023b), for each object, we randomly selected two different frames from the same video clip to form a training pair. We apply the object mask on one frame to obtain the background-free object as the input reference object. For the other frame, we use the bounding box of the object as the grounding information and use this frame as the training ground truth. For single-image data, we use LVIS (Gupta et al., 2019), a well-known dataset for fine-grained large vocabulary instance segmentation, including 118,287 images from 1,203 categories. For each sample, we select only the object instances with top-10 largest bounding box area as training data.
@@ 4.2 Evaluation Metrics
We calculate the CLIP-I (Radford et al., 2021) score and DINO (Caron et al., 2021) score to evaluate the identity preservation performance of the subjects and use CLIP-T (Radford et al., 2021) score to evaluate the text alignment performance of the generated image. For evaluation of the model’s grounding ability, we use $AP_{50}$ based on a pretrained YOLOv8 (Jocher et al., 2023) object detection model.
Figure 4: Visual comparison with existing methods on DreamBench objects for the single-subject customization task. Please zoom in to see the details.
Figure 5: Multi-concept customization on DreamBench objects. Please zoom in to see the details.
4.3 Single Subject Customization
We compare our work with existing state-of-the-art works on DreamBench (Ruiz et al., 2023) for the customization of a single subject. In this experiment, we use the bounding box of the subject in the ground-truth image as the input layout. The qualitative and quantitative results are shown in Fig. 4 and Table 1, respectively. Overall, our method shows significantly better performance in layout alignment, reference object identity preservation, and background text alignment. Existing encoder-based subject-driven text-to-image customization methods BLIP-Diffusion (Li et al., 2024), ELITE (Wei et al., 2023), $\lambda$-eclipse (Patel et al., 2024) and MLLM-based method Kosmos-G (Pan et al., 2023) fail to maintain accurate identity of the reference objects lack the ability for precise layout control. AnyDoor (Chen et al., 2023b) is designed for image composition on a given background and lacks the ability of text-to-image generation. Although previous grounded text-to-image generation methods like GLIGEN (Li et al., 2023) are able to achieve layout control, it cannot preserve the identity of the subjects. CustomNet (Yuan et al., 2023) achieves flexible pose control. However, it highly relies on the pretrained model Zero123 (Liu et al., 2023), limiting the resolution of its generated image to be $256\times 256$. Moreover, there can be obvious artifacts around the boundary of the generated subject.
Table 1: Comparison with existing methods on Dreambench.
@@ 4.3 Single Subject Customization
SD V1.4 [(Rombach et al., 2022)] 0.3122 0.8413 0.6587
BLIP-Diffusion [(Li et al., 2024)] 0.2824 0.8894 0.7625
ELITE [(Wei et al., 2023)] 0.2461 0.8936 0.7557
Kosmos-G [(Pan et al., 2023)] 0.2864 0.8452 0.6933
lambda-eclipse [(Patel et al., 2024)] 0.2767 0.8973 0.7934
AnyDoor [(Chen et al., 2023b)] 0.2430 0.9062 0.7928
GLIGEN [(Li et al., 2023)] 0.2898 0.8520 0.6890
CustomNet [(Yuan et al., 2023)] 0.2815 0.9090 0.7526
@@
As an interesting observation, we find that previous non-grounding based customization methods are inclined to generate objects that are very large and in the center of the image, which gains benefit in CLIP-I score and DINO score during evaluation. However, in real-world scenarios, users may expect to flexibly control the size of the subject in the generated images. They may choose to generate larger background with broader textual information, where, in such cases, non-grounding customization methods cannot generate the desired result. The visual results in Fig. 4 demonstrate that our results achieves better identity preservation performance with accurate layout-alignment. We encourage the readers to view more visualizations in the Appendix.
4.4 Multi-Subject Customization and Multi-Entity Background Generation
With our proposed masked cross-attention module, our model seamlessly supports the customization of multiple subjects. Fig. 5 shows the qualitative results of the task where we customize multiple subjects and generate the image by grounding multiple text entities in the background. It can be observed that when inputting multiple subjects such as a bag and a boot, along with the layout of the background text entities such as the mountain and the lake, our model successfully generates the subjects and background with an accurate layout-alignment of each visual concept. The generated subjects preserve the their identity well. The overall generated image is well text-aligned and artifact free. Moreover, in several cases, even when the bounding boxes of the foreground objects have a large overlap with the background text entities, the model can distinguish subject-driven foreground generation from text-driven background generation, effectively avoiding the context blending.
Figure 6: Visaul results of reference-guided image generation with complex layout and text entities as conditions on COCO validation set. Note that LayoutDiffusion (Zheng et al., 2023) is only conducted on COCO dataset with filtered annotations, so some of its results are not available.
4.5 Customization with Complex Layout and Text Entities
We evaluate our model’s performance the COCO validation set for the task of generating customized images with complex layout and text entities as guidance. Quantitative and qualitative results are shown in Table 2 and Fig. 6, respectively. For each testing image, we use the largest object as the reference object (i.e., the subject), and the remaining text entities as background entities. To quantify the model’s grounding ability, we adopt YOLOv8 (Jocher et al., 2023) as the object detection method. Results show that even if we input complex layouts and text entities to the model, our model can still generate high-quality scenes with precise layout alignment of all the objects and regions, and accurate identity preservation of the reference object, while preserving the text-alignment. Compared with previous layout-to-image generation methods, our model has a competitive accuracy in grounding the visual concepts and remarkable improvement on identity preservation.
Table 2: Quantitative results on MS-COCO validation set for the task of customized image generation with complex layout as guidance. In this setting, we finetune our model on COCO training set, and compare with previous methods that only train on COCO.
CLIP-T ↑ CLIP-I ↑ DINO-I ↑ $AP_{50}$
LAMA[ (Li et al., 2021)] 0.2507 0.8441 0.7330 18.20
LayoutDiffusion[ (Zheng et al., 2023)] 0.2738 0.8655 0.8033 27.40
UniControl[ (Qin et al., 2023)] 0.3143 0.8425 0.7598 4.53
GLIGEN[ (Li et al., 2023)] 0.2899 0.8688 0.7792 27.50
Ours 0.2946 0.9078 0.8560 31.10
4.6 Ablation Study
We conduct the ablation study to validate the effectiveness of our proposed components: the masked cross-attention module and the grounding module. Table 4 and Table 4 present the quantitative results on DreamBench and COCO, respectively. We observe that both the grounding module and the masked cross-attention module play a vital role in the model’s grounding ability and prevent the information leakage of the reference object. Benefiting from these two modules, the model shows stronger ability of identity preservation, text alignment and grounded generation.
@@ 4.6 Ablation Study
CLIP-T ↑ CLIP-I ↑ DINO-I ↑ $AP_{50}$
w/o Grounding Module 0.2796 0.8605 0.7740 22.00
w/o Masked Cross-Attention 0.2884 0.8707 0.7970 28.50
Full 0.2946 0.9078 0.8560 31.10
Table 5: User Study based on DreamBench: In the questions, the user is presented side-by-side comparisons of our generated image and another image randomly chosen from one of the baselines. The results in the table show user preference percentage.
Ours CustomNet Ours AnyDoor Ours GLIGEN
Identity 60.78 39.22 59.31 40.69 72.81 27.19
Grounding 56.86 43.14 64.21 35.79 58.25 41.75
Text Alignment 51.96 48.03 73.52 26.47 55.34 44.66
Overall Quality 54.41 45.58 62.25 37.74 58.74 41.26
4.7 User Study
In Table 5, we show the user study results comparing our model with existing models (Chen et al., 2023b; Yuan et al., 2023; Li et al., 2023) on DreamBench. Specifically, given the same input, we generate results with each model. Then we ask the users to make side-by-side comparison of our result and a randomly chosen result from the baselines regarding identity preservation, text alignment, grounding ability, and overall image quality. We collect the user responses using Amazon Mechanical Turk. Results show that participants have significantly higher preference on our method. We show details about user study in the Appendix Sec. D.
5 Conclusion and Future Work
@@
Although our model successfully generates customized images with layout control, there are still several limitations. First, the model’s performance can be limited by the base model. We can address this by using a stronger base model. Second, the design of reusing the masked cross-attention layer for each subject could still be time-consuming during inference. This can be addressed by developing a parallel generation structure for multiple subjects. We leave this direction in future work.
Appendix
Appendix A Preliminary
Our model is based on Stable Diffusion v1.4 Rombach et al. (2022), a Latent Diffusion model (LDM) that applies the diffusion process in a latent space. Specifically, an input image $x$ is encoded into the latent space using a pretrained autoencoder $z=\mathcal{E}(x),\hat{x}=\mathcal{D}(z)$ (with an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$). Then the denoising process is achieved by training a denoiser $\epsilon_{\theta}\left(z_{t},t,f_{c}\right)$ that predicts the added noise following:
$\min_{\theta}E_{z_{0},\epsilon\sim\mathcal{N}(0,1),t\sim\mathrm{U}(1,T)}\left\|\epsilon-\varepsilon_{\theta}\left(z_{t},t,f_{c}\right)\right\|_{2}^{2},$ (8)
where $f_{c}$ is the embedding of the condition (such as a prompt) and $z_{t}$ is the latent noise at timestamp $t$.
Appendix B Training/Inference Details
Our model is trained on 4 NVIDIA A100 GPUs for 100k steps with a batch size of 14 and a learning rate of $5\times 10^{-5}$. During training, we randomly drop reference image embedding and text embedding both at the rate of 10%. We decently rank the area of the boxes per images, and set the max number of grounding boxes to be 10 with the largest areas. During inference, we set classifier-free guidance(CFG) (Ho & Salimans, 2022) as 3.
Appendix C Details about Data Collection
For each reference image, we use the segmentation mask to mask out the background and get the background-free reference object. In inference stage, we use SAM (Kirillov et al., 2023) to get the mask of the reference object, and get the background-free reference object.
Appendix D Details about User Study
Our user study is based on DreamBench, with full 30 objects and 25 prompts. We randomly generated layouts, and use them in the generation. In the user study, given the layout, the reference object, the text prompt, the result of our method and a random-selected baseline method, we request the user to answer the following four questions:
(1) Which generated image do you think that its object is more similar to the input object? Choose between Option A and B.
(2) Which generated image do you think that its object is most likely to be at the right position as the input layout? Choose between Option A and B.
(3) Which generated image do you think is most likely to match the text description? Choose between Option A and B.
(4) Which image do you think has better image quality? Choose between Option A and B.
We received more than 1200 votes from over 530 users. In the experiment, we randomly shuffle the order of baselines to improve the confidence of the user study.
Appendix E Additional Qualititive Results
Here we show more qualitative results. In Fig. 7 we show results on DreamBench and in Fig. 8 and Fig. 9 we show more results about complex background background evaluation on coco validation set.
Figure 7: More visual results of our model.
Appendix F Social Impact
GroundingBooth provides a flexible method for users to precisely customize the layout of both foreground and background objects based on user-provided reference subjects and text descriptions without any test-time finetuning. The support for the generation of multi-subjects provides a useful tool for users to generate images using their desired layout. Users can optionally choose reference objects or simple text inputs to generate their desired image, which significantly expand the flexibility in controllable and customized text-to-image generation.
Figure 8: More results on complex scene generation on COCO validation set.
Figure 9: More results on complex scene generation on COCO validation set.